# Every nonflat conformal minimal surface is homotopic to a proper one

### v1

@@
{quoting}
[leftmargin=5mm] Abstract  Given an open Riemann surface $M$, we prove that every nonflat conformal minimal immersion $M\to\mathbb{R}^{n}$ ($n\geq 3$) is homotopic through nonflat conformal minimal immersions $M\to\mathbb{R}^{n}$ to a proper one. If $n\geq 5$, it may be chosen in addition injective, hence a proper conformal minimal embedding. Prescribing its flux, as a consequence, every nonflat conformal minimal immersion $M\to\mathbb{R}^{n}$ is homotopic to the real part of a proper holomorphic null embedding $M\to\mathbb{C}^{n}$. We also obtain a result for a more general family of holomorphic immersions from an open Riemann surface into $\mathbb{C}^{n}$ directed by Oka cones in $\mathbb{C}^{n}$.
@@ 1.  Introduction
We also have the following corollary, which provides information on the topology of the space of proper nonflat conformal minimal immersions.
@@
Problem 1.4.
@@ 2.  Notation and preliminaries
We begin the preparatory section by establishing notation and recalling some basic definitions. Denote by =$\mathbb{N}=\{1,2,\ldots\}$ the set of natural numbers, by +=$\mathbb{Z}_{+}=\mathbb{N}\cup\{0\}$ the set of non-negative integers, and assume that $n\in\mathbb{N}$, $n\geq 3$. Let ||$\cdot$|| ⋅ | denote the Euclidean norm in $\mathbb{R}^{n}$, and $||\cdot||_{\infty}$ the infinity norm in $\mathbb{R}^{n}$. Let $K$ be a compact topological space and let $f:K\to\mathbb{R}^{n}$ be a continuous map. We write $\|f\|_{K}$ for the maximum norm of $f$ on $K$. Given maps $f,g\colon K\to\mathbb{R}^{n}$, we say that $f$ approximates $g$ on $K$, and write $f\approx g$, if $||f-g||_{K}<\varepsilon$ for some $\varepsilon>0$ which is sufficiently small according to the argument. If not specified otherwise, throughout the paper, we assume that surfaces are connected.
Let $X,\ Y$ be complex manifolds and $M\subset X$ a subset of $X$. By $\mathscr{C}^{r}(M,Y)$, $r\in\mathbb{N}\cup\{0\}$, we denote the space of $r$-times continuously differentiable maps $M\to Y$, and by $\mathscr{O}(M,Y)$ the space of holomorphic maps from some unspecified neighbourhood of $M$ in $X$ to $Y$. If, in addition, we are given a closed subset $K\subset X$, then we write $\mathscr{A}(K,Y)$ for the space of continuous maps $K\to Y$ which are holomorphic in the interior $\mathring{K}\subset K$ of $K$, i.e., $\mathscr{A}(K,Y)=\mathscr{C}^{0}(K,Y)\cap\mathscr{O}(\mathring{K},Y)$. In the case when $Y=\mathbb{C}$, we omit writing the latter in the corresponding spaces of functions, and therefore consider the spaces $\mathscr{C}^{r}(M)$, $\mathscr{O}(M)$, and $\mathscr{A}(K)$.
@@
Minimal surfaces appear as the real and the imaginary parts of holomorphic null curves, that is, holomorphic immersions $z=(z_{1},\dots,z_{n})\colon M\to\mathbb{C}^{n}$ ($n\geq 3$) from an open Riemann surface $M$ into $\mathbb{C}^{n}$ that satisfy $\sum_{i=1}^{n}z_{i}^{2}=0$. Conversely, given a simply connected domain $D\subset M$ and a conformal minimal immersion $u\colon M\to\mathbb{R}^{n}$, the restriction $u|_{D}\colon D\to\mathbb{R}^{n}$ is the real part of a holomorphic null curve $z\colon D\to\mathbb{C}^{n}$. Moreover, $u\colon M\to\mathbb{R}^{n}$ globally corresponds to the real part of a holomorphic null curve if and only if it holds that $\int_{C}d^{c}u=0$ for any closed curve $C\subset M$. In particular, if the latter is satisfied, i.e., the conjugate differential $d^{c}u=2\Im(\partial{u})$ is an exact $1$-form on $M$, then $u$ admits a harmonic conjugate $v$, and $z=u+\imath v\colon M\to\mathbb{C}^{n}$ is a holomorphic null curve (here, $\imath^{2}=-1$ denotes the imaginary unit).
@@ 3.  Nearby conformal minimal surfaces are isotopic
This section is devoted to the main new ingredient, Lemma 3.1. It says that given a nonflat conformal minimal immersion $u\colon K\to\mathbb{R}^{n}$ defined on a smoothly bounded compact domain $K$, if there is another nonflat conformal minimal immersion $\tilde{u}\colon K\to\mathbb{R}^{n}$ uniformly close to $u$ on $K$, then they are homotopic through nonflat conformal minimal immersions. This somehow upgrades the property of being close to being homotopic, which is one of the mayor features we employ in the proof of the main result. As explained in Section 6, it enables us to improve some other results concerning approximation of minimal surfaces defined on compact bordered Riemann surfaces, and we expect to establish further applications.
@@
We choose the number $\varepsilon_{1}>0$ so small, such that if $\tilde{f}\in\mathscr{O}(K,{\bf A}_{*})$ is as above and the holomorphic $1$-forms $f_{t}\theta$ have vanishing real periods for $t\in[0,1]$, i.e., $\Re\mathscr{P}(f_{t}\theta)=0$ (see (3)), then
@@ 4.  The noncritical and critical cases in the induction
    $||\hat{u}_{1}(x)||_{\infty}>\delta$ for all $x\in L\setminus\mathring{K}$ and $||\hat{u}_{1}(x)||_{\infty}>1/\varepsilon$ for all $x\in bL$.
@@
Pick a number $\varepsilon_{1}\in(0,\varepsilon/3)$. Note that $u_{1}\in\mathrm{CMI}_{*}(K,\mathbb{R}^{n})$ maps $bK$ into $\mathbb{R}^{n}\setminus\{0\}$ (4.1). Hence, by [6, Lemma 3.11.1, Theorem 3.4.1], we may approximate $u_{1}$ uniformly on $K$ by $u_{2}\in\mathrm{CMI}_{*}(L,\mathbb{R}^{n})$ such that $||u_{2}(x)||_{\infty}>\delta$ for all $x\in L\setminus\mathring{K}$, $||u_{2}(x)||_{\infty}>1/\varepsilon$ for all $x\in bL$, $\mathrm{Flux}(u_{2})=\mathrm{Flux}(u_{1})$, and $u_{2}$ is injective if $n\geq 5$. Further, Lemma 3.1, applied to $(u,\tilde{u})=(u_{1},u_{2})$, provides a homotopy $u_{p}\in\mathrm{CMI}_{*}(K,\mathbb{R}^{n})$, $p\in[0,2]$, with the following properties:
@@
    $u_{2}\in\mathrm{CMI}_{*}(L,\mathbb{R}^{n})$, $||u_{2}(x)||_{\infty}>\delta$ for all $x\in L\setminus\mathring{K}$, $||u_{2}(x)||_{\infty}>1/\varepsilon$ for all $x\in bL$, $\mathrm{Flux}(u_{2})=\mathrm{Flux}(u_{1})$, $u_{2}$ is injective if $n\geq 5$;
@@
for some $r\in(1/2,1)$. Provided that approximation of $f_{p,\zeta}$ by $g_{p,\zeta}$ is sufficiently close on $K$, the implicit function theorem furnishes us with a continuous map
@@
To finish the proof, we shall reparametrize the homotopy $\tilde{u}_{p}\in\mathrm{CMI}_{*}(L,\mathbb{R}^{n})$, $p\in[0,2]$, to obtain a homotopy $\hat{u}_{p}\in\mathrm{CMI}_{*}(L,\mathbb{R}^{n})$, $p\in[0,1]$, meeting the conclusions of the proposition. There exists a small number $\tau>0$ close to 00, such that
@@
we obtain a homotopy of nonflat conformal minimal immersions $L\to\mathbb{R}^{n}$ satisfying the conclusions of the proposition. Indeed, $\hat{u}_{0}=\tilde{u}_{0}=u_{0}$, $\hat{u}_{1}=\tilde{u}_{2}=u_{2}$, so $||\hat{u}_{1}(x)||_{\infty}=||u_{2}(x)||_{\infty}>\delta$ for all $x\in L\setminus\mathring{K}$, $||\hat{u}_{1}(x)||_{\infty}>1/\varepsilon$ for all $x\in bL$, $\mathrm{Flux}(\hat{u}_{1})=\mathrm{Flux}(u_{2})=\mathrm{Flux}(u_{1})$, and $\hat{u}_{1}$ is injective if $n\geq 5$. For the uniform approximation of $u_{p}$ by $\hat{u}_{p}$ on $K$ we argue as follows.
@@
  • $S=K\cup E$ is an admissible set in $M$ such that $E$ closes inside $K$ to a closed Jordan curve $C$ such that $E=\overline{C\setminus K}$.
  • $\Omega\subset\mathbb{R}^{n}$ is a domain containing $0\in\mathbb{R}^{n}$, $P=[0,1]$, $x_{0}\in K\cap E$.
@@
  • $u_{p}\in\mathrm{CMI}_{*}(K,\mathbb{R}^{n})$ is a continuous family of nonflat conformal minimal immersions and $f_{p}=2\partial{u}_{p}/\theta\colon K\to{\bf A}_{*}$ is a continuous family of nonflat holomorphic maps for every $p\in P$.
@@
(Recall that $f_{p}^{t}=f_{p}$ (4.5) holds on $S$.) We shall change $\sigma_{p}$ on $I_{1}\cup I_{2}$ in such a way that the modified maps $\tilde{\sigma}_{p}^{t}$ will satisfy $\Re\int_{C}\tilde{\sigma}_{p}^{t}=0$ for all $(p,t)\in P\times[t_{0},1]$, for some value $t_{0}\in(0,1)$. At the same time, $\sigma_{p}$ will remain unchanged on $[0,3]\setminus(I_{1}\cup I_{2})$.
@@
[…]
@@
In the second step, we modify $\sigma_{p}$ on $I_{2}$ to insure that the integrals (4.7) equal 00. By assumption on $f_{p}$, $\sigma_{p}$ is nonflat on $I_{2}$ for all $p\in P$. Therefore, [3, Lemma 3.6] allows us to embed $\sigma_{p}|_{I_{2}}$ as the core $\tau_{p,0}=\sigma_{p}$ of a period dominating spray of paths
@@
    $||u_{1}^{i}(x)||_{\infty}>i-1$ for all $x\in{M_{i}\setminus\mathring{M}_{i-1}}$.
@@
    $20<\varepsilon_{i}<\varepsilon_{i-1}/$ 2, and every harmonic map $v\colon M_{i}\to\mathbb{R}^{n}$ such that $||v-u_{t}^{i}||_{M_{i}}<2\varepsilon_{i}$ for some $t\in[0,1]$ is a nonflat immersion, and is in addition injective if $n\geq 5$ and the inequality holds for $t=1$.
@@
Case 2: The domain $A_{i}=\{c_{i-1}<\rho<c_{i}\}$ contains a critical point $x$ of $\rho$. By assumption, $x$ is a unique critical point of $\rho$ on $A_{i}$ and it is a Morse point with Morse index either 00 or $1$. We consider the following subcases.
Subcase 2.1: The Morse index of $\rho$ at $x$ equals 00. Take a small smoothly bounded disc $\Delta\subset A_{i}$ around $x$ in $A_{i}$. Extend $u_{t}^{i-1}\in\mathrm{CMI}_{*}(M_{i-1},\mathbb{R}^{n})$, $t\in[0,1]$, to a disjoint union $S_{i}=M_{i-1}\cup\Delta$ by defining any homotopy of nonflat conformal minimal immersions $u_{t}^{i-1}\in\mathrm{CMI}_{*}(\Delta,\mathbb{R}^{n})$, $t\in[0,1]$, such that $u_{0}^{i-1}=u|_{\Delta}$ and $||u_{1}^{i-1}(q)||_{\infty}>i-2$ for all $q\in\Delta$. The set $S_{i}=M_{i-1}\cup\Delta\subset M_{i}$ is admissible in $M_{i}$ and has the same topology as $M_{i}$, therefore $u_{t}^{i}\in\mathrm{CMI}_{*}(M_{i},\mathbb{R}^{n})$ is obtained by applying Case 1.
@@
(5.1) $\begin{cases}u_{0}^{i-1}=u|_{S_{i}},\\ ||u_{1}^{i-1}(q)||_{\infty}>i-2\quad\text{for all }q\in E,\\ \mathrm{Flux}(u_{1}^{i-1})(C)=\mathfrak{p}(C).\end{cases}$
@@
$\begin{cases}\tilde{u}_{0}^{i-1}=u|_{V_{i}},\\ ||\tilde{u}_{1}^{i-1}(q)||_{\infty}>i-2\quad\text{for all }q\in V_{i}\setminus\mathring{M}_{i-1},\\ \mathrm{Flux}(\tilde{u}_{1}^{i-1})=\mathfrak{p}\quad\text{on }H_{1}(S_{i},\mathbb{Z}).\end{cases}$

### v2

@@
  • Abstract  Given an open Riemann surface $M$, we prove that every nonflat conformal minimal immersion $M\to\mathbb{R}^{n}$ ($n\geq 3$) is homotopic through nonflat conformal minimal immersions $M\to\mathbb{R}^{n}$ to a proper one. If $n\geq 5$, it may be chosen in addition injective, hence a proper conformal minimal embedding. Prescribing its flux, as a consequence, every nonflat conformal minimal immersion $M\to\mathbb{R}^{n}$ is homotopic to the real part of a proper holomorphic null embedding $M\to\mathbb{C}^{n}$. We also obtain a result for a more general family of holomorphic immersions from an open Riemann surface into $\mathbb{C}^{n}$ directed by Oka cones in $\mathbb{C}^{n}$.
@@ 1. Introduction
We also have the following two corollaries, which provide information on the topology of the space of proper nonflat conformal minimal immersions.
@@
Corollary 1.4.
Let $M$ be an open Riemann surface and $H_{1}(M,\mathbb{Z})\cong\mathbb{Z}^{l}$ with $l\in\mathbb{Z}_{+}\cup\{\infty\}$. Then the space $\mathrm{CMI}_{*}^{p}(M,\mathbb{R}^{3})$ of proper nonflat conformal minimal immersions $M\to\mathbb{R}^{3}$ has at least $2^{l}$ connected components.
The second statement is obtained in view of [3, Proposition 8.4] and the discussion therein.
Problem 1.5.
@@ 2. Notation and preliminaries
We begin the preparatory section by establishing notation and recalling some basic definitions. Denote by $\mathbb{N}=\{1,2,\ldots\}$ the set of natural numbers, by $\mathbb{Z}_{+}=\mathbb{N}\cup\{0\}$ the set of non-negative integers, and assume that $n\in\mathbb{N}$, $n\geq 3$. Let $|\cdot|$ denote the Euclidean norm in $\mathbb{R}^{n}$, and $||\cdot||_{\infty}$ the infinity norm in $\mathbb{R}^{n}$. Let $K$ be a compact topological space and let $f:K\to\mathbb{R}^{n}$ be a continuous map. We write $\|f\|_{K}$ for the maximum norm of $f$ on $K$. Given maps $f,g\colon K\to\mathbb{R}^{n}$, we say that $f$ approximates $g$ on $K$, and write $f\approx g$, if $f-g||_{K}<\varepsilon$ for some $\varepsilon>0$ which is sufficiently small according to the argument. If not specified otherwise, throughout the paper, we assume that surfaces are connected.
Let X,YX,\ Y be complex manifolds and $M\subset X$ a subset of $X$. By $\mathscr{C}^{r}(M,Y)$, $r\in\mathbb{N}\cup\{0\}$, we denote the space of $r$-times continuously differentiable maps $M\to Y$, and by $\mathscr{O}(M,Y)$ the space of holomorphic maps from some unspecified neighbourhood of $M$ in $X$ to $Y$. If, in addition, we are given a closed subset $K\subset X$, then we write $\mathscr{A}(K,Y)$ for the space of continuous maps $K\to Y$ which are holomorphic in the interior $\mathring{K}\subset K$ of $K$, i.e., $\mathscr{A}(K,Y)=\mathscr{C}^{0}(K,Y)\cap\mathscr{O}(\mathring{K},Y)$. In the case when $Y=\mathbb{C}$, we omit writing the latter in the corresponding spaces of functions, and therefore consider the spaces $\mathscr{C}^{r}(M)$, $\mathscr{O}(M)$, and $\mathscr{A}(K)$.
@@
Minimal surfaces appear as the real and the imaginary parts of holomorphic null curves, that is, holomorphic immersions $z=(z_{1},\dots,z_{n})\colon M\to\mathbb{C}^{n}$ ($n\geq 3$) from an open Riemann surface $M$ into $\mathbb{C}^{n}$ that satisfy $\sum_{i=1}^{n}(dz_{i})^{2}=0$. Conversely, given a simply connected domain $D\subset M$ and a conformal minimal immersion $u\colon M\to\mathbb{R}^{n}$, the restriction $u|_{D}\colon D\to\mathbb{R}^{n}$ is the real part of a holomorphic null curve $z\colon D\to\mathbb{C}^{n}$. Moreover, $u\colon M\to\mathbb{R}^{n}$ globally corresponds to the real part of a holomorphic null curve if and only if it holds that $\int_{C}d^{c}u=0$ for any closed curve $C\subset M$. In particular, if the latter is satisfied, i.e., the conjugate differential $d^{c}u=2\Im(\partial{u})$ is an exact $1$-form on $M$, then $u$ admits a harmonic conjugate $v$, and $z=u+\imath v\colon M\to\mathbb{C}^{n}$ is a holomorphic null curve (here, $\imath^{2}=-1$ denotes the imaginary unit).
@@ 3. Nearby conformal minimal surfaces are isotopic
This section is devoted to the main new ingredient, Lemma 3.1. It says that given a nonflat conformal minimal immersion $u\colon K\to\mathbb{R}^{n}$ defined on a smoothly bounded compact domain $K$, if there is another nonflat conformal minimal immersion $\tilde{u}\colon K\to\mathbb{R}^{n}$ uniformly close to $u$ on $K$, then they are homotopic through nonflat conformal minimal immersions. This somehow upgrades the property of being close to being homotopic, which is one of the major features we employ in the proof of the main result. As explained in Section 6, it enables us to improve some other results concerning approximation of minimal surfaces defined on compact bordered Riemann surfaces, and we expect to establish further applications.
@@
We choose the number $\varepsilon_{1}>0$ so small, such that if $\tilde{f}\in\mathscr{O}(K,{\bf A}_{*})$ is as above and the holomorphic $1$-forms $f_{t}\theta$ have vanishing real periods for $t\in[0,1]$, i.e., $\Re\mathscr{P}(f_{t})=0$ (see (3)), then
@@ 4. The noncritical and critical cases in the induction
    $\hat{u}_{1}(x)||_{\infty}>\delta$ for all $x\in L\setminus\mathring{K}$ and $\hat{u}_{1}(x)||_{\infty}>1/\varepsilon$ for all $x\in bL$.
@@
Pick a number $\varepsilon_{1}\in(0,\varepsilon/3)$. Note that $u_{1}\in\mathrm{CMI}_{*}(K,\mathbb{R}^{n})$ maps $bK$ into $\mathbb{R}^{n}\setminus\{0\}$ (4.1). Hence, by [6, Lemma 3.11.1, Theorem 3.4.1], we may approximate $u_{1}$ uniformly on $K$ by $u_{2}\in\mathrm{CMI}_{*}(L,\mathbb{R}^{n})$ such that $u_{2}(x)||_{\infty}>\delta$ for all $x\in L\setminus\mathring{K}$, $u_{2}(x)||_{\infty}>1/\varepsilon$ for all $x\in bL$, $\mathrm{Flux}(u_{2})=\mathrm{Flux}(u_{1})$, and $u_{2}$ is injective if $n\geq 5$. Further, Lemma 3.1, applied to $(u,\tilde{u})=(u_{1},u_{2})$, provides a homotopy $u_{p}\in\mathrm{CMI}_{*}(K,\mathbb{R}^{n})$, $p\in[0,2]$, with the following properties:
@@
    $u_{2}\in\mathrm{CMI}_{*}(L,\mathbb{R}^{n})$, $u_{2}(x)||_{\infty}>\delta$ for all $x\in L\setminus\mathring{K}$, $u_{2}(x)||_{\infty}>1/\varepsilon$ for all $x\in bL$, $\mathrm{Flux}(u_{2})=\mathrm{Flux}(u_{1})$, $u_{2}$ is injective if $n\geq 5$;
@@
for some $r\in(1/2,1)$, and such that it satisfies $g_{p,0}=f_{p}$ for $p\in\{0,2\}$. Provided that approximation of $f_{p,\zeta}$ by $g_{p,\zeta}$ is sufficiently close on $K$, the implicit function theorem furnishes us with a continuous map
@@
To finish the proof, we shall reparametrize the homotopy $\tilde{u}_{p}\in\mathrm{CMI}_{*}(L,\mathbb{R}^{n})$, $p\in[0,2]$, to obtain a homotopy $\hat{u}_{p}\in\mathrm{CMI}_{*}(L,\mathbb{R}^{n})$, $p\in[0,1]$, meeting the conclusions of the proposition. There exists a small number $\tau>0$ close to 0, such that
@@
[…]
@@
  • $S=K\cup E$ is an admissible set in $M$ such that $E$ closes inside $K$ to a closed Jordan curve $C$ with $E=\overline{C\setminus K}$ and $K\cap E=\{x_{0},x_{1}\}$.
@@
  • $u_{p}\in\mathrm{CMI}_{*}(K,\mathbb{R}^{n})$ is a continuous family of nonflat conformal minimal immersions and $f_{p}=2\partial{u}_{p}/\theta\colon K\to{\bf A}_{*}$ is a continuous family of nonflat holomorphic maps for every $p\in P=[0,1]$.
  • $\Omega\subset\mathbb{R}^{n}$ is a domain containing $0\in\mathbb{R}^{n}$ and the point $u_{1}(x_{1})-u_{1}(x_{0})$.
@@
(Recall that $f_{p}^{t}=f_{p}$ (4.5) holds on $S$.) We shall change $\sigma_{p}$ on $I_{1}\cup I_{2}$ in such a way that the modified maps $\tilde{\sigma}_{p}^{t}$ will satisfy $\Re\int_{C}\tilde{\sigma}_{p}^{t}=0$ for all $(p,t)\in P\times[t_{0},1]$, for some value $t_{0}\in(0,1)$. At the same time, $\sigma_{p}$ will remain unchanged on $0,3]\setminus(I_{1}\cup I_{2})$.
@@
(4.7) $\left|\Re\int_{0}^{3}\sigma_{p}^{t}(s)ds\right|<\varepsilon,\quad(p,t)\in P\times[t_{0},1],$
@@
In the second step, we modify $\sigma_{p}$ on $I_{2}$ to insure that the integrals (4.7) equal 0. By assumption on $f_{p}$, $\sigma_{p}$ is nonflat on $I_{2}$ for all $p\in P$. Therefore, [3, Lemma 3.6] allows us to embed $\sigma_{p}|_{I_{2}}$ as the core $\tau_{p,0}=\sigma_{p}$ of a period dominating spray of paths
@@
    $u_{1}^{i}(x)||_{\infty}>i-1$ for all $x\in{M_{i}\setminus\mathring{M}_{i-1}}$.
@@
    $0<\varepsilon_{i}<\varepsilon_{i-1}/2$, and every harmonic map $v\colon M_{i}\to\mathbb{R}^{n}$ such that $v-u_{t}^{i}||_{M_{i}}<2\varepsilon_{i}$ for some $t\in[0,1]$ is a nonflat immersion, and is in addition injective if $n\geq 5$ and the inequality holds for $t=1$.
@@
Case 2: The domain $A_{i}=\{c_{i-1}<\rho<c_{i}\}$ contains a critical point $x$ of $\rho$. By assumption, $x$ is a unique critical point of $\rho$ on $A_{i}$ and it is a Morse point with Morse index either 0 or $1$. We consider the following subcases.
Subcase 2.1: The Morse index of $\rho$ at $x$ equals 0. Take a small smoothly bounded disc $\Delta\subset A_{i}$ around $x$ in $A_{i}$. Extend $u_{t}^{i-1}\in\mathrm{CMI}_{*}(M_{i-1},\mathbb{R}^{n})$, $t\in[0,1]$, to a disjoint union $S_{i}=M_{i-1}\cup\Delta$ by defining any homotopy of nonflat conformal minimal immersions $u_{t}^{i-1}\in\mathrm{CMI}_{*}(\Delta,\mathbb{R}^{n})$, $t\in[0,1]$, such that $u_{0}^{i-1}=u|_{\Delta}$ and $u_{1}^{i-1}(q)||_{\infty}>i-2$ for all $q\in\Delta$. The set $S_{i}=M_{i-1}\cup\Delta\subset M_{i}$ is admissible in $M_{i}$ and has the same topology as $M_{i}$, therefore $u_{t}^{i}\in\mathrm{CMI}_{*}(M_{i},\mathbb{R}^{n})$ is obtained by applying Case 1.
@@
[…]
@@
[…]